Find the sum of vector A and C
A=64i + 64j
C=64i - 128j
1
Expert's answer
2020-02-27T10:09:13-0500
"\bold{A+C}=(64\bold{i}+64\bold{j})+(64\bold{i}-128\bold{j})"

"\bold{A+C}=(64+64)\bold{i}+(64-128)\bold{j}"


"\bold{A+C}=128\bold{i}-64\bold{j}"